Hello i have a problem with the current problem. The problem is this:
Jen invested $6000, part at 10% and the rest at 5%. If the annual income from both investments was the same, find her total annual income from both investments.
The choices include:
A)$200
B)$250
C)$300
D)$350
E)$400

Answer by Earlsdon(6294)

If x = the amount Jen invested at 10%, then the remainder (6000-x) was invested at 5%. The earnings on these two amounts are equal, so...
0.1x = 0.05(6000-x) Simplify and solve for x.
0.1x = 300-0.05x Add 0.05x to both sides.
0.15x = 300 Divide both sides by 0.15
x = 2000 and 6000-x = 4000
The interest earned on the $2000 at 10% = 0.1(2000) = $200
The interest earned on the $4000 at 5% = 0.05(4000) = $200
The total interest earned = $400